EE 488 HW 15 For the three-phase transformers shown, the individual windings are rated: primary H winding 7200v secondary
Answer #1

Given:

VP-L-L=VP-L-N =13.2KV

Rated Individual Windings Voltage

H = 7.2KV

X = 0.48KV

M = 1.4KV

Assumption:

Let us assume that the Voltage rating of the H windings are given as V L-N

Proof:

a.

The turn ratio for Primary and H winding is VP-L-N  / V S- L-N = 13.2/7.2 = 1.833, Here the phase difference is 30o because of different windings

Now for the Primary and X windings, the turns ration is VP-L-N  / V S- L-N = 13.2 / 0.48 = 27.5, Here the Phase shift is 0o between primary and secondary because of similar windings.